Write the equation of the line that is perpendicular to the line y = 2x + 2 and passes through the point (6, 3).
Answer by math-vortex(472)

Hi, there--

The Problem:
Write the equation of the line that is perpendicular to the line y = 2x + 2 and passes through 
the point (6, 3).

The Solution:
We need to find the slope of the perpendicular line. Then we can find an equation for the line 
using point-slope form.

The original equation y = 2x + 2 is in slope-intercept form (y=mx+b) so we can read the slope 
right off the equation. The slope m is 2.

The slope of the perpendicular line will always have the slope -1/m. In this case, the slope of 
the perpendicular line is -1/m=-1/2.

The point-slope form for a line is y-y[1] = m(x-x[1]), where m is the slope and (x[1],y[1]) is a point on the line. 

The point (x[1],y[1]) is (6,3), so the equation is
y - (3) = (-1/2)(x - 6)

You can simplify this a bit if we want to, but it's not essential.
y-3 = (-1/2)x+3
y = (-1/2)x+6
.
Feel free to email if you have questions about the solution.

Ms.Figgy
math.in.the.vortex@gmail.com